OpenStudy (anonymous):

Simplify: (-xy^0x^6a^5)^-3

OpenStudy (anonymous):

\[n^a\cdot n^b=n^{a+b}\] \[\left( n^a \right)^b=n^{ab}\] \[n^0=1, n\neq 0\]

OpenStudy (anonymous):

Those are the three laws you'll need

OpenStudy (anonymous):

I'd start by getting rid of the y term, then multiply the -3 exponent in

OpenStudy (anonymous):

Okay so I got here: (-xy^0x^6a^5)^-3 (-x^7y^0a^5)-3 Is that right so far?

OpenStudy (anonymous):

yep so far so good

OpenStudy (anonymous):

wait hang on that -3 is still an exponent, you're missing a ^

OpenStudy (anonymous):

probably just a typo though right? :)

OpenStudy (anonymous):

Yup, thanks :) Okay so..

OpenStudy (anonymous):

Now I have to distribute the -3 to everything

OpenStudy (anonymous):

make your life easier first; what's y^0?

OpenStudy (anonymous):

It's 1

OpenStudy (anonymous):

right, so you can essentially just drop that term, since it's just multiplying stuff by 1 inside the brackets

OpenStudy (anonymous):

Okay so: (-1)^-3, (1)^-3, (a^5)^-3 right?

OpenStudy (anonymous):

\[\left( -x^7a^5 \right)^{-3}=\left( \left( -1 \right)^{-3}x^{7\cdot -3}a^{5\cdot -3} \right)\]

OpenStudy (anonymous):

that's multiplication in the exponents, it's not super clear on my screen maybe i should have used brackets

OpenStudy (anonymous):

I am so confused. What did I do wrong?

OpenStudy (anonymous):

i'm not sure what you were doing when you said (-1)^-3, (1)^-3 but (a^5)^-3 was correct, you just also needed to account for (x^7)^-3

OpenStudy (anonymous):

Okay so: 3x^14a^-10

OpenStudy (anonymous):

how did you get 3, 14, -10?

OpenStudy (anonymous):

You said above: ((-1)^-3x^-7*-2a^5*-2)

OpenStudy (anonymous):

So i simplified it.

OpenStudy (anonymous):

\[-1^{-3}=-1\]

OpenStudy (anonymous):

\[\left( x^7 \right)^{-3}=x^{7(-3)}\]

OpenStudy (anonymous):

\[\left( a^5 \right)^{-3}=a^{5(-3)}\]

OpenStudy (anonymous):

does that make sense to you?

OpenStudy (anonymous):

-1x^-14a^-10 ?

OpenStudy (anonymous):

what's 7 times -3?

OpenStudy (anonymous):

and 5 times -3

OpenStudy (anonymous):

that loos like it says -2

OpenStudy (anonymous):

ah the font on this site sucks XD

OpenStudy (anonymous):

-1x^-15a^-21

OpenStudy (anonymous):

lol \[\Large -x^{-21}a^{-15}\]

OpenStudy (anonymous):

i figured it out, you do \Large in the equation :D

OpenStudy (anonymous):

are you confused now or did that help

OpenStudy (anonymous):

cause we can talk about exponents some more in a simpler approach

OpenStudy (anonymous):

oops i switched them :P

OpenStudy (anonymous):

ya it's k :) you had the right idea i just confused you with that tiny exponent font

OpenStudy (anonymous):

Okay so I cant let them have negitives. So: 1/x^21a^15 right? \[\frac{ 1 }{ x^21a^15 }\]

OpenStudy (anonymous):

put your exponents in curly brackets

OpenStudy (anonymous):

*x^21a^15

OpenStudy (anonymous):

and remember your term is negative

OpenStudy (anonymous):

\[\frac{ -1 }{ x^{21}a^{15} }\]

OpenStudy (anonymous):

Yup! Got it! Thanks so much!

OpenStudy (anonymous):

my pleasure